# CREST: An Efficient Conjointly-trained Spike-driven Framework for Event-based Object Detection Exploiting Spatiotemporal Dynamics

### v1

@@ Abstract
Event-based cameras feature high temporal resolution, wide dynamic range, and low power consumption, which is ideal for high-speed and low-light object detection. Spiking neural networks (SNNs) are promising for event-based object recognition and detection due to their spiking nature but lack efficient training methods, leading to gradient vanishing and high computational complexity, especially in deep SNNs. Additionally, existing SNN frameworks often fail to effectively handle multi-scale spatiotemporal features, leading to increased data redundancy and reduced accuracy. To address these issues, we propose CREST, a novel conjointly-trained spike-driven framework to exploit spatiotemporal dynamics in event-based object detection. We introduce the conjoint learning rule to accelerate SNN learning and alleviate gradient vanishing. It also supports dual operation modes for efficient and flexible implementation on different hardware types. Additionally, CREST features a fully spike-driven framework with a multi-scale spatiotemporal event integrator (MESTOR) and a spatiotemporal-IoU (ST-IoU) loss. Our approach achieves superior object recognition & detection performance and up to 100$\times$ energy efficiency compared with state-of-the-art SNN algorithms on three datasets, providing an efficient solution for event-based object detection algorithms suitable for SNN hardware implementation.
@@ Introduction
The FSN works as follows. Firstly, the FSN $n$ in layer $l$ integrates the input spike $\delta_{m}^{l-1}$ to the membrane potential $U_{n}^{l}$ (Eq. 1). $N^{l-1}$, $K^{l-1}$, $d^{l}(t)$, $w_{mn}^{l}$ represent the number of input neurons, time-window length, spike coefficient at time $t$, and connection weights respectively. If $u_{n}^{l}(t)$ exceeds the threshold $U_{th}^{l+1}(t)$, a spike is fired (Eq. 2), and $u_{n}^{l}(t)$ decays by $U_{th}^{l+1}(t)$ (Eq. 3). The neuron dynamics parameters $U_{th}^{l+1}(t)$ and $d^{l}(t)$ can be used to model various activation functions.
$U_{n}^{l}$ $=\sum_{m=1}^{N^{l-1}}\sum_{t=1}^{K^{l}}d^{l}(t)\delta_{m}^{l-1}(t)w_{mn}^{l}$ (1)
$\delta_{n}^{l}(t)$ =$=\begin{cases}1,&u_{n}^{l}(t)\geq U_{th}^{l+1}(t)\\ 0,&\text{otherwise }\end{cases}$ (2)
$u_{n}^{l}(t+1)$ $=U_{n}^{l}-\sum_{t}U_{th}^{l+1}(t)\delta_{n}^{l}(t)$ (3)
@@
4:         $\delta_{m}^{l-1}(t)$leftarrow $Spiking(U_{m}^{l-1},\alpha^{l},K^{l})$triangleright Eq.(2)
@@
6:     $x_{n}^{l}=U_{n}^{l}$leftarrow $Integrate(\delta_{m}^{l-1}(t),\alpha^{l},K^{l},w_{mn}^{l})$triangleright Eq.(1)
@@
8:     $x_{q}^{l-1}\leftarrow Clip(x_{m}^{l-1},\alpha^{l},K^{l})$triangleright Eq.(10)
9:     $x_{p}^{l-1}\leftarrow Round(x_{q}^{l-1},\alpha^{l},K^{l})$triangleright Eq.(11)
@@
14:$\frac{\partial x_{n}^{l}}{\partial x_{m}^{l-1}}=\frac{\partial x_{n}^{l}}{\partial x_{p}^{l-1}}\frac{\partial x_{p}^{l-1}}{\partial x_{q}^{l-1}}=\frac{\partial x_{n}^{l}}{\partial x_{p}^{l-1}}$
@@
ANN-based object detection mostly adopts ReLU and leaky-ReLU (Glorot, Bordes, and Bengio 2011; Maas et al. 2013). We implement ReLU in this paper and the detailed reasons are illustrated in Supplementary Material B. Consider the connection between FSN $n$ in layer $l$ and FSN $p$ in layer $l+1$. FSN $n$ integrates the membrane potential $U_{n}^{l}$ (Eq. 6) based on Eq. 3. The output spikes from FSN $n$ to FSN $p$ are integrated unweighted (Eq. 7) based on Eq. 1.
$U_{n}^{l}\!\!=\!u_{n}^{l}(t\!+\!1)\!+\!\!\!\sum_{t}\!U^{l+1}_{th}(t)\delta^{l}_{n}(t)\!\!=\!\!\!\sum_{t=1}^{K^{l+1}\!\!-\!1}\!\!U_{th}^{l+1}(t)\delta_{n}^{l}(t)\!+\!u_{n}^{l}(K)$ (6)
$U_{p}^{l+1}=\sum_{t=1}^{K^{l+1}}d^{l+1}(t)\delta_{n}^{l}(t)$ (7)
@@
We can observe that FSN $n$ encodes membrane potentials into spike trains, which FSN $p$ then decodes back into continuous values, effectively replacing the MAC in ANNs with AC in SNNs. In this process, $U_{n}^{l}$ and $U_{p}^{l+1}$ emulate the input and output of ReLU. To preserve ReLU’s non-linearity while controlling the output range, ReLU is bounded by $\alpha$. This is formulated as three cases: ❶ when $U_{n}^{l}\!\leq\!0$, $U_{p}^{l+1}\!=\!0$; ❷ when $0\!<\!U_{n}^{l}\!\leq\!\alpha$, $U_{p}^{l+1}\!=\!U_{n}^{l}$; and ❸ when $U_{n}^{l}\!>\!\alpha$, $U_{p}^{l+1}\!\!=\!\alpha$. Case ❷ is optimized by minimizing the error with proper $d^{l+1}(t)$ and $U_{th}^{l+1}(t)$ under the distribution of $U_{n}^{l}$ (Eq. 8), with $K^{l+1}$ pre-set by accuracy and firing rate needs.
$\textbf{min}\;\varepsilon(U_{th}^{l+1}(t),d^{l+1}(t);K^{l+1})=\mathbb{E}_{u_{n}^{l}(t)}|U_{n}^{l}-U_{p}^{l+1}|$ (8)
A possible setting is $U_{th}^{l+1}(t)\!\!=\!d^{l+1}(t)\!\!=\!\alpha^{l+1}2^{-t}$. For case ❶, when $u_{n}^{l}(t)\!\leq\!0$, FSN $n$ generates no spikes ($U_{p}^{l+1}\!\!=\!0$). For case ❷, when $u_{n}^{l}(t)$ is an integer multiple of $\alpha^{l+1}2^{-K^{l+1}}$, $U_{p}^{l+1}\!=\!U_{n}^{l}$. For case ❸, when $u_{n}^{l}(t)\!\!\geq\!\alpha^{l+1}(1\!-\!2^{-K^{l+1}})$, the spike train is encoded as $K$-bit 1 with $U_{p}^{l+1}\!=\!\alpha^{l+1}(1\!-\!2^{-K^{l+1}})$. The error between $U_{p}^{l+1}$ and $\alpha^{l+1}$ decreases with larger $K^{l+1}$. Otherwise, $U_{n}^{l}$ rounds down to the nearest level with a tolerance $\alpha^{l+1}2^{-K^{l+1}}$. Notably, both earlier spikes and higher spiking frequency represent larger values (more significant information), as reflected by larger coefficients ($d(t)$) and accumulated values, which improves encoding efficiency.
@@
FSN-BP requires complex and iterative spatiotemporal gradient BP with high computation complexity and slow convergence, which can exacerbate gradient vanishing in deep SNNs. Inspired by the coarse-to-fire-processing strategy of FSN-ReLU, we design a surrogate neural network with discrete-level activation values (DL-Net) to approximate spike train patterns in Fig. 1(c).
The temporal iterative ACs in Eq. 1 are replaced by MACs (Eq. 9), where $x_{n}^{l}$ emulates $U_{n}^{l}$. Then, to emulate the boundary conditions of the 3 cases, $x_{n}^{l}$ is clipped to $x_{q}^{l}$ (Eq. 10) with $X^{l}\!=\!\alpha^{l+1}(1-2^{-\!K^{l+1}}\!)$. Further, $x_{q}^{l}$ is discretized to mimic the FSN spiking process with $X^{l}_{\text{min}}\!\!=\alpha^{l+1}2^{-K^{l+1}}$ (Eq. 11). This approach converts the original spatiotemporal accumulation into spatial accumulation and discretization.
$x^{l}_{n}=\sum_{m=1}^{N^{l-1}}x_{m}^{l-1}w_{mn}^{l}$ (9)
$x_{q}^{l}=\text{clip}(x^{l}_{n},0,X^{l})=\begin{cases}0,&x^{l}_{n}<0\\ x^{l}_{n},&0\leq x^{l}_{n}\leq X^{l}\\ X^{l},&\text{otherwise}\end{cases}$ (10)
$x_{p}^{l}=\text{round}(\frac{x_{q}^{l}}{X^{l}_{\text{min}}})\times X^{l}_{\text{min}}$ (11)
@@
YOLE (Cannici et al. 2019) VoxelGrid ANN - 0.927 - 1 0.33G 3.04
@@
Asynet (Messikommer et al. 2020) VoxelGrid ANN - 0.931 - 0.067 0.32G 0.19
HybridSNN (Kugele et al. 2021) HIST SNN - 0.770 - - - 1.13
@@
VGG-11 (Cordone, Miramond, and Thierion 2022) VoxelCube SNN 9.23M 0.924 5 0.147 0.61G 12.37
DenseNet121-16 (Fan et al. 2024) VoxelCube SNN 1.76M 0.937 3 0.147 0.06G* 0.73
@@
Table 1: Comparison between CREST and the state-of-the-art methods on NCAR dataset.
@@
For a continuous event-based period $\mathcal{I}$, the raw stream $E\!\!=\!\!{e_{i}(x_{i},y_{i},t_{i},p_{i})}_{i\in N}$ with total $N$ bipolar events $p_{i}$ at pixel point $(x,y)$ exhibits high sparsity and temporal redundancy. These characteristics impede feature learning and increase computational overhead. Moreover, the raw stream contains many redundant background and noise events, which can disrupt the learning process. Existing methods mostly split $\mathcal{I}(E)$ into equal bins and integrate the events in each bin to obtain a spatiotemporal dense representation (Eq. 13) (Su et al. 2023; Fan et al. 2024). The bin length affects the spatiotemporal feature preservation, leading to a trade-off between both scales. We aim to integrate the input feature from multi-scales (spatiotemporal scale $\mathcal{ST}$, spatial scale $\mathcal{S}$ and temporal scale $\mathcal{T}$) for better representation.
@@
An FSN-based convolution layer is constructed which applies $N$ fixed-value filters of size $3\times 3$ to the $N$ channel time bins. In detail, the convolution results of each time bin and each filter are accumulated to the membrane potential $\widehat{u}$ in each time step (Eq. 15). Then, based on Eq.2 and 3, each neuron generates a spike train (Eq. 16). With a given time window $K$, if the number of spikes in one spike train is greater than $K-1$, this neuron (pixel) is considered space-time continuous and we set the value of this pixel as 1 (we retain this pixel). Otherwise, we abandon this pixel (Eq. 17).
@@
Asynet (Messikommer et al. 2020) VoxelGrid ANN YOLO 11.4M 0.129 - - 0.100 2.09G 0.96
@@
VC-Mobile (Cordone, Miramond, and Thierion 2022) VoxelCube SNN SSD 12.64M 0.174 - 5 0.294 4.34G 76.22
@@
We adopt the commonly used NCAR dataset for object recognition (Sironi et al. 2018), Gen1 and PKU-Vidar-DVS dataset for object detection (de Tournemire et al. 2020; Li et al. 2022a). The details of these datasets are shown in Supplementary Material D.
@@
Set $K=5$ and $\alpha=3$ for middle layers, with $\alpha=1$ for the input encoding layer in FSN-SNN models. All models are trained on a Tesla V100 GPU using the AdamW optimizer and StepLR scheduler. For recognition, training spans 100 epochs with a batch size of 64, an initial learning rate of $31e-$ 3, and a weight decay of $21e-$ 2. For detection, training also spans 100 epochs with a batch size of 32, an initial learning rate of $31e-$ 3, and a weight decay of $45e-$ 4. Data augmentation techniques include mosaic, random rotation, and flipping.
@@
CREST-A 0.279 0.185 5.06×5.06$\times$5.06 ×
CREST-B 0.358 0.173 5.08×5.08$\times$5.08 ×
CREST 0.360 0.167 4.96×4.96$\times$4.96 ×
@@
MESTOR-A 0.873 0.164 0.859 0.187 0.871 0.151
MESTOR-B 0.938 0.185 0.922 0.199 0.937 0.154
MESTOR 0.949 0.165 0.940 0.176 0.952 0.147
Table 6: Ablation Study: comparison between the conventional representation and MESTOR on NCARS dataset.
@@
The conjoint learning rule reduces training time per epoch by 5×5$\times$5 × and enhances mAP performance by reducing errors caused by surrogate gradient backpropagation across multiple dimensions. MESTOR and ST-IoU further boost mAP and lower firing rates by fully leveraging the spatiotemporal feature of the event data.

### v3

@@ Abstract
Event-based cameras feature high temporal resolution, wide dynamic range, and low power consumption, which are ideal for high-speed and low-light object detection. Spiking neural networks (SNNs) are promising for event-based object recognition and detection due to their spiking nature but lack efficient training methods, leading to gradient vanishing and high computational complexity, especially in deep SNNs. Additionally, existing SNN frameworks often fail to effectively handle multi-scale spatiotemporal features, leading to increased data redundancy and reduced accuracy. To address these issues, we propose CREST, a novel conjointly-trained spike-driven framework to exploit spatiotemporal dynamics in event-based object detection. We introduce the conjoint learning rule to accelerate SNN learning and alleviate gradient vanishing. It also supports dual operation modes for efficient and flexible implementation on different hardware types. Additionally, CREST features a fully spike-driven framework with a multi-scale spatiotemporal event integrator (MESTOR) and a spatiotemporal-IoU (ST-IoU) loss. Our approach achieves superior object recognition & detection performance and up to 100$\times$ energy efficiency compared with state-of-the-art SNN algorithms on three datasets, providing an efficient solution for event-based object detection algorithms suitable for SNN hardware implementation.
@@ Introduction
The FSN works as follows. Firstly, the FSN $n$ in layer $l$ integrates the input spike $\delta_{m}^{l-1}$ to the membrane potential $U_{n}^{l}$ (Eq. 1). $N^{l-1}$, $K^{l-1}$, $d^{l-1}(t)$, $w_{mn}^{l}$ represent the number of input neurons, time-window length, spike coefficient at time $t$, and connection weights respectively. If $u_{n}^{l}(t)$ exceeds the threshold $U_{th}^{l}(t)$, a spike is fired (Eq. 2), and $u_{n}^{l}(t)$ decays by $U_{th}^{l}(t)$ (Eq. 3). The neuron dynamics parameters $U_{th}^{l}(t)$ and $d^{l}(t)$ can be used to model various activation functions.
$U_{n}^{l}$ $=\sum_{m=1}^{N^{l-1}}\sum_{t=1}^{K^{l-1}}d^{l-1}(t)\delta_{m}^{l-1}(t)w_{mn}^{l}$ (1)
$\delta_{n}^{l}(t)$ =$=\begin{cases}1,&u_{n}^{l}(t)\geq U_{th}^{l}(t)\\ 0,&\text{otherwise }\end{cases}$ (2)
$u_{n}^{l}(t+1)$ $=U_{n}^{l}-\sum_{t}U_{th}^{l}(t)\delta_{n}^{l}(t)$ (3)
@@
4:         $\delta_{m}^{l-1}(t)$leftarrow $Spiking(u_{m}^{l-1}(t),\alpha^{l-1},K^{l-1})$triangleright Eq.(2)
@@
6:     $x_{n}^{l}=U_{n}^{l}$leftarrow $Integ.(\delta_{m}^{l-1}(t),\alpha^{l-1},K^{l-1},w_{mn}^{l})$triangleright Eq.(1)
@@
8:     $x_{q}^{l-1}\leftarrow Clip(x_{m}^{l-1},\alpha^{l-1},K^{l-1})$triangleright Eq.(9)
9:     $x_{p}^{l-1}\leftarrow Round(x_{q}^{l-1},\alpha^{l-1},K^{l-1})$triangleright Eq.(10)
@@
14:$\frac{\partial x_{n}^{l}}{\partial x_{m}^{l-1}}=\frac{\partial x_{n}^{l}}{\partial x_{p}^{l-1}}\frac{\partial x_{p}^{l-1}}{\partial x_{q}^{l-1}}\approx\frac{\partial x_{n}^{l}}{\partial x_{p}^{l-1}}$
@@
ANN-based object detection mostly adopts ReLU and leaky-ReLU (Glorot, Bordes, and Bengio 2011; Maas et al. 2013). We implement ReLU in this paper and the detailed reasons are illustrated in Supplementary Material B. Based on Eq. 3, rewrite the membrane potential $U_{m}^{l-1}$ as Eq. 6. Meanwhile, denote the unweighted integration of output spikes from FSN $m$ as Eq. 7 based on Eq. 1.
$U_{m}^{l-1}=\sum_{t=1}^{K^{l-1}}U_{th}^{l-1}(t)\delta_{m}^{l-1}(t)+u_{m}^{l-1}(K^{l-1})$ (6)
$U_{p}^{l-1}=\sum_{t=1}^{K^{l-1}}d^{l-1}(t)\delta_{m}^{l-1}(t)$ (7)
@@
With the help of FSN parameters, the membrane potential is encoded into a spike train and then decoded into a continuous value. In this process, $U_{m}^{l-1}$ and $U_{p}^{l-1}$ emulate the input and output of ReLU, and the MAC in ANNs are replaced by the ACs of the corresponding weight at spiking time steps in SNNs. To preserve non-linearity while controlling the output range, ReLU is bounded by $\alpha$. This is formulated as three cases: ❶ when $U_{m}^{l-1}\!\leq\!0$, $U_{p}^{l-1}\!=\!0$; ❷ when $0\!<\!U_{m}^{l-1}\!\leq\!\alpha$, $U_{p}^{l-1}\!=\!U_{m}^{l-1}$; and ❸ when $U_{m}^{l-1}\!>\!\alpha$, $U_{p}^{l-1}\!\!=\!\alpha$. Case ❷ is optimized by minimizing the error with proper FSN parameters under the distribution of $U_{m}^{l-1}$ (Eq. 8), with $K^{l-1}$ pre-set by the requirements of accuracy and firing rate.
$\textbf{min}\;\varepsilon(U_{th}^{l-1}(t),d^{l-1}(t);K^{l-1})=\mathbb{E}_{u_{m}(t)}|U_{m}^{l-1}-U_{p}^{l-1}|$ (8)
A possible setting is $U_{th}^{l-1}(t)\!\!=\!d^{l-1}(t)\!\!=\!\alpha^{l-1}2^{-t}$. For case ❶, when $U_{m}^{l-1}\!\leq\!0$, FSN $m$ generates no spikes ($U_{p}^{l-1}\!\!=\!0$). For case ❷, when $U_{m}^{l-1}$ is an integer multiple of $\alpha^{l-1}2^{-K^{l-1}}$, $U_{p}^{l-1}\!=\!U_{m}^{l-1}$. Otherwise, $U_{m}^{l-1}$ rounds down to the nearest level with a tolerance $\alpha^{l-1}2^{-K^{l-1}}$. For case ❸, $U_{m}^{l-1}(t)\!\!\geq\!\alpha^{l-1}2^{-t}$, the spike train is encoded as $K$-bit 1 with $U_{p}^{l-1}\!=\!\alpha^{l-1}(1\!-\!2^{-K^{l-1}})$. The error between $U_{p}^{l-1}$ and $U_{m}^{l-1}$ decreases when $K^{l-1}$ increases. Noted that both the earlier spikes and the higher spiking frequency represent larger values (more significant information), as reflected by larger coefficients ($d(t)$) and accumulated values, which improves encoding efficiency.
@@
FSN-BP requires complex and iterative spatiotemporal gradient BP with high computation complexity, which can exacerbate gradient vanishing in deep SNNs. Inspired by the coarse-to-fire-processing strategy of FSN-ReLU, we design a surrogate neural network with discrete-level activation values (DL-Net) to approximate spike train patterns in Fig. 1(c).
To emulate the boundary condition of the 3 cases, $x_{m}^{l-1}$ is clipped to $x_{q}^{l-1}$ (Eq. 9) with $X^{l-1}\!=\!\alpha^{l-1}(1-2^{-\!K^{l-1}}\!)$. Further, $x_{q}^{l-1}$ is discretized to mimic the FSN spiking process with $X^{l-1}_{\text{min}}\!\!=\alpha^{l-1}2^{-K^{l-1}}$ (Eq. 10). Finally, the temporal iterative ACs in Eq. 1 are replaced by MACs (Eq. 11), where $x_{n}^{l}$ emulates $U_{n}^{l}$. The connecting weights $w_{pn}^{l}$ in DL-Net are shared with those in the FSN-SNN.
$x_{q}^{l-1}=\text{clip}(x^{l-1}_{m})=\begin{cases}0,&x^{l-1}_{m}<0\\ x^{l-1}_{m},&0\leq x^{l-1}_{m}\leq X^{l-1}\\ X^{l-1},&\text{otherwise}\end{cases}$ (9)
$x_{p}^{l-1}=\text{round}(\frac{x_{q}^{l-1}}{X^{l-1}_{\text{min}}})\times X^{l-1}_{\text{min}}$ (10)
$x^{l}_{n}=\sum_{p=1}^{N^{l-1}}x_{p}^{l-1}w_{pn}^{l}$ (11)
@@
YOLE (Cannici et al. 2019) VoxelGrid ANN - 0.927 - 1 0.16G 0.75
@@
Asynet (Messikommer et al. 2020) VoxelGrid ANN - 0.944 - 0.067 0.16G 0.05
HybridSNN (Kugele et al. 2021) HIST SNN - 0.770 - - - -
@@
VGG-11 (Cordone, Miramond, and Thierion 2022) VoxelCube SNN 9.23M 0.924 5 0.120 0.61G 12.68
DenseNet121-16 (Fan et al. 2024) VoxelCube SNN 1.76M 0.937 5 0.147 0.06G* 1.22
@@
Table 1: Comparison between CREST and the state-of-the-art methods on NCARs dataset.
@@
For a continuous event-based period $\mathcal{I}$, the raw stream $E\!\!=\!\!{e_{i}(x_{i},y_{i},t_{i},p_{i})}_{i\in M}$ with total $M$ bipolar events $p_{i}$ at pixel point $(x_{i},y_{i})$ exhibits high sparsity and temporal redundancy. These characteristics impede feature learning and increase computational overhead. Moreover, the raw stream contains many redundant background and noise events, which can disrupt the learning process. Existing methods mostly split $\mathcal{I}(E)$ into equal bins and integrate the events in each bin to obtain a spatiotemporal dense representation (Eq. 13) (Su et al. 2023; Fan et al. 2024). The bin length affects the spatiotemporal feature preservation, leading to a trade-off between both scales. We aim to integrate the input feature from multi-scales (spatiotemporal scale $\mathcal{ST}$, spatial scale $\mathcal{S}$ and temporal scale $\mathcal{T}$) for better representation.
@@
An FSN-based convolution layer is constructed which applies fixed-value filters of size $3\times 3$ to time bin $\mathbf{\mathcal{B}}_{t}$. In detail, the convolution results of each time bin and each filter are accumulated to the membrane potential $\widehat{u}$ in each time step (Eq. 15). Then, based on Eq.2 and 3, each neuron generates a spike train (Eq. 16). With a given time window $K$, if the number of spikes in one spike train is greater than $K-1$, this neuron (pixel) is considered space-time continuous and we set the value of this pixel as 1 (we retain this pixel). Otherwise, we abandon this pixel (Eq. 17).
@@
Asynet (Messikommer et al. 2020) VoxelGrid ANN YOLO 11.4M 0.129 - - 0.100 1.05G 0.48
@@
VC-Mobile (Cordone, Miramond, and Thierion 2022) VoxelCube SNN SSD 24.26M 0.147 - 5 0.294 4.34G 76.22
@@
We adopt the commonly used NCARs dataset for object recognition (Sironi et al. 2018), Gen1 and PKU-Vidar-DVS dataset for object detection (de Tournemire et al. 2020; Li et al. 2022a). The details of these datasets are shown in Supplementary Material D.
@@
Set $K=5$ and $\alpha=3$ for middle layers, with $\alpha=1$ for the input encoding layer in FSN-SNN models. All models are trained on a Tesla V100 GPU using the AdamW optimizer and StepLR scheduler. For recognition, training spans 100 epochs with a batch size of 64, an initial learning rate of $31e-$ 3, and a weight decay of $21e-$ 2. For detection, training also spans 100 epochs with a batch size of 32, an initial learning rate of $31e-$ 3, and a weight decay of $45e-$ 4. Data augmentation techniques include mosaic, random rotation, crop and flipping.
@@
CREST-A 0.279 0.185 3.42×3.42$\times$3.42 ×
CREST-B 0.358 0.173 3.43×3.43$\times$3.43 ×
CREST 0.360 0.167 3.41×3.41$\times$3.41 ×
@@
MESTOR-A 0.873 0.164 0.871 0.151 0.859 0.187
MESTOR-B 0.938 0.185 0.937 0.154 0.922 0.199
MESTOR 0.949 0.165 0.952 0.146 0.940 0.176
Table 6: Ablation study: effectiveness of MESTOR’s components on NCARs dataset.
@@
The conjoint learning rule reduces training time per epoch by 3.4×3.4$\times$3.4 × and enhances mAP performance by reducing errors caused by surrogate gradient backpropagation across multiple dimensions. MESTOR and ST-IoU further boost mAP and lower firing rates by fully leveraging the spatiotemporal feature of the event data.